# AppReferences

### v1

@@ 5.1 Data and Study Design
The final dataset, obtained by merging the previously mentioned data sources and cleaning the data set to avoid missing data, comprises 384 counties, 270 of which were designed as attainment and 114 as non-attainment. In Appendix, Figure F.1 maps this final analysis data set.
@@ 5.2 Results
We apply the proposed CASBAH model to the 384 counties in the eastern US described above, including all covariates —- census and meteorological variables —- in the weights of the post-treatment variable mixture, while for the outcome model we use the linear model in (10). The model identifies the three strata: the dissociative stratum with 115 counties ($30\%$ of the total counties analyzed), the associative positive stratum with 48 counties ($12.5\%$), and the associative negative stratum with 221 counties ($57.5\%$). According to the definition of the three strata and as visualized in the image on the left in Figure 2, the dissociative stratum—identified with the color yellow—is composed of counties where the NAAQS revision does not substantially affect the level of pm2.5 , in fact, the expected value of $\mathbb{E}[P_{i}(1)-P_{i}(0)|S_{i}(1)=S_{i}(0)]$ for the counties allocated to this strata has a median close to zero and the first and third quantiles of $-0.29\mu g/m^{3}$ and $0.07\mu g/m^{3}$. The associative negative stratum—identified with the color green—is made up of counties where the implementation of environmental plans significantly decreases the levels of pm2.5 . Specifically, in these counties, the NAAQS revision reduced by $0.9\mu g/m^{3}$ the median levels pm2.5 , with first and third quantiles of $-1.46\mu g/m^{3}$ and $-0.38\mu g/m^{3}$. The associative positive stratum—identified with the color red—is made up of counties where the revision of NAAQS was associated with increases in pm2.5 levels by $0.8\mu g/m^{3}$ in the median. However, this latter stratum has the highest uncertainty for $\mathbb{E}[P_{i}(1)-P_{i}(0)|S_{i}(1)\prec S_{i}(0)]$, with values of the first and third quantiles equal to $0.13\mu g/m^{3}$ and $1.34\mu g/m^{3}$, respectively.
@@
The corresponding distributions of the expected dissociative/associative effects are reported in the right image in Figure 2 and show the effect of the attainment or non-attainment designations on the mortality rate conditional to the three strata—i.e., conditional to the heterogeneity in the causal effects of the NAAQS revision on the level of pollution. The EDE and the EAE+ have a similar increase in the median $2.7{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ of the mortality rate when environmental plans are applied. The means of these two effects are $0.8{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ and $1.3{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$, respectively. For both principal causal effects, the interquartile interval includes the zero value. This means that in counties where the NAAQS revision does not affect the level pm2.5 or where this level increases, the mortality rate is also not affected. The EAE-—in green color on the right image of Figure 2—assumes negative values, indicating that the implementation of environmental plans, in the counties where the regulations affect the level of pm2.5 , significantly reducing it, also reduces the median age adjusted mortality rate. Specifically, the mortality rate decreases by $11.9{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ in the median and with an interquartile interval of $[-15.2{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}},-9.1{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}]$.
@@
In addition, it is our interest to characterize the three different strata (Figure F.3 in the Supplementary Materials visualizes the average of observed covariates within each stratum—reported as colored lines—and compares them with the average of covariates between the full 384 counties). The associative positive stratum and the dissociative stratum are composed of rural counties with a smaller population density, where the population has lower income with respect to the mean of the overall counties, and with a small employment rate. In contrast, in countries of the associative positive stratum, the percentage of Hispanics is lower than the average, in favor of a higher percentage of whites and Asians, while the dissociative stratum is characterized by a higher percentage of Hispanics and a smaller community of black people, with a higher level of education. Furthermore, the associative positive stratum is distinguished by a high smoke rate. In contrast, the associative negative stratum identifies urban counties with a high population density. This population is characterized by higher income, but also a higher percentage of poor people and lower levels of education. The meteorological variable, such as the averages of daily temperatures, the relative average of humidity, and the dew point, appears to play an important role in the distinction of strata—that is, in the characterization of the different effects of the implementation of environmental plans on the level of pm2.5. We refer to the Supplementary Materials for more details about the probability of each county to belong in each of the three strata.
6 Conclusion
In this paper, we proposed the Confounders-Aware SHared Atoms Bayesian Mixture Model (CASBAH). CASBAH uses the dependent Dirichlet process to create an highly flexible and adaptable distribution for the potential post-treatment variables, taking into account both the confounders and the treatment variable. This approach extends the recent BNP model for causal inference proposed by Zorzetto et al. (2024) to the principal stratification settings. CASBAH also provides an estimate of causal effects for a set of innovative estimates of principal causal effects.
The proposed model improves nonparametric modeling for principal stratification, e.g., the model proposed by Schwartz et al. (2011), by offering a more flexible approach to handling post-treatment variables. CASBAH does so by integrating the confounders into the weights of an infinite Bayesian mixture model, improving control over measured confounding, and enhancing the precision of causal effect estimates. Additionally, by utilizing shared random kernel atoms, the model facilitates the identification of associative or dissociative strata without predetermined criteria, effectively learning from the data across different treatment levels.
CASBAH dynamically identifies the main strata and imputes missing post-treatment variables and outcomes without the need to set predefined thresholds on the post-treatment values, as is done in most of the principal stratification literature in the context of continuous post-treatment. Moreover, the model quantifies the uncertainty in membership in the principal strata and estimates the principal causal effects. Theorem 1 provides a useful factorization of the expected value for each potential outcome, while Theorem 2 illustrates how the probability of a unit belonging to a dissociative stratum is non-zero and depends on its observed covariates.
@@ 6 Conclusion
The performance of the proposed model is evaluated in a battery of Monte Carlo simulations where the performance of CASBAH is compared and contrasted with the performance of previous BNP modeling for principal stratification proposed by Schwartz et al. (2011). Furthermore, the proposed model is used to assess the effectiveness of a previous revision of NAAQS in reducing levels of pm2.5and, in turn, reducing the mortality rate in treated counties. Specifically, we find significant effects on the reduction of mortality rate in areas where pollution was effectively reduced by the non-attainment designation (associative negative effects). In contrast, those areas where pollution was not measurably affected (dissociative effect) or where it actually increased (associative positive effects) did not show any significant effect on the mortality rate.

### v2

@@ 5.1 Data and Study Design
The final dataset, obtained by merging the previously mentioned data sources and cleaning the data set to avoid missing data, comprises 384 counties, 270 of which were designed as attainment and 114 as non-attainment. In the Supplementary Material, Figure F.1 maps this final analysis data set.
@@ 5.2 Results
We apply our proposed model to the 384 counties in the eastern US described above, including all covariates, census and meteorological variables, in the weights of the post-treatment variable mixture, while for the outcome model we use the linear model in (10). The model identifies the three strata: the dissociative stratum with 115 counties ($30\%$ of the total counties analyzed), the associative positive stratum with 48 counties ($12.5\%$), and the associative negative stratum with 221 counties ($57.5\%$). According to the definition of the three strata and as visualized in the image on the left in Figure 2, the dissociative stratum, identified with the color yellow, is composed of counties where the NAAQS revision does not substantially affect the level of pm2.5 , in fact, the expected value of $\mathbb{E}[P_{i}(1)-P_{i}(0)|S_{i}^{(1)}=S_{i}^{(0)}]$ for the counties allocated to this strata has a median close to zero and the first and third quantiles of $-0.29\mu g/m^{3}$ and $0.07\mu g/m^{3}$. The associative negative stratum, identified with the color green, is made up of counties where the implementation of environmental plans significantly decreases the levels of pm2.5 . Specifically, in these counties, the NAAQS revision reduced by $0.9\mu g/m^{3}$ the median levels pm2.5 , with first and third quantiles of $-1.46\mu g/m^{3}$ and $-0.38\mu g/m^{3}$. The associative positive stratum, identified with the color red, is made up of counties where the revision of NAAQS was associated with increases in pm2.5 levels by $0.8\mu g/m^{3}$ in the median. However, this latter stratum has the highest uncertainty for $\mathbb{E}[P_{i}(1)-P_{i}(0)|S_{i}^{(1)}\prec S_{i}^{(0)}]$, with values of the first and third quantiles equal to $0.13\mu g/m^{3}$ and $1.34\mu g/m^{3}$, respectively.
@@
The corresponding distributions of the expected dissociative/associative effects are reported in the right image in Figure 2 and show the effect of the attainment or non-attainment designations on the mortality rate conditional to the three strata, i.e., conditional to the heterogeneity in the causal effects of the NAAQS revision on the level of pollution. The EDE and the EAE+ have a similar increase in the median $2.7{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ of the mortality rate when environmental plans are applied. The means of these two effects are $0.8{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ and $1.3{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$, respectively. For both principal causal effects, the interquartile interval includes the zero value. This means that in counties where the NAAQS revision does not affect the level pm2.5 or where this level increases, the mortality rate is also not affected. The EAE-, in green color on the right image of Figure 2, assumes negative values, indicating that the implementation of environmental plans, in the counties where the regulations affect the level of pm2.5 , significantly reducing it, also reduces the median age adjusted mortality rate. Specifically, the mortality rate decreases by $11.9{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ in the median and with an interquartile interval of $[-15.2{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}},-9.1{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}]$.
@@
In addition, it is our interest to characterize the three different strata (Figure F3 in the Supplementary Materials visualizes the average of observed covariates within each stratum, reported as colored lines, and compares them with the average of covariates between the full 384 counties). The associative positive stratum and the dissociative stratum are composed of rural counties with a smaller population density, where the population has lower income with respect to the mean of the overall counties, and with a small employment rate. In contrast, in countries of the associative positive stratum, the percentage of Hispanics is lower than the average, in favor of a higher percentage of whites and Asians, while the dissociative stratum is characterized by a higher percentage of Hispanics and a smaller community of black people, with a higher level of education. Furthermore, the associative positive stratum is distinguished by a high smoke rate. In contrast, the associative negative stratum identifies urban counties with a high population density. This population is characterized by higher income, but also a higher percentage of poor people and lower levels of education. The meteorological variable, such as the averages of daily temperatures, the relative average of humidity, and the dew point, appears to play an important role in the distinction of strata, that is, in the characterization of the different effects of the implementation of environmental plans on the level of pm2.5. We refer to the Supplementary Materials for more details about the probability of each county to belong in each of the three strata.
6 Discussion
In this paper, we proposed the confounders-aware shared atoms Bayesian mixture model to address critical challenges within the principal stratification framework. Our approach leverages the dependent Dirichlet process to define a highly flexible and adaptable model for the potential post-treatment variables. Our proposed idea of allowing the distribution of potential post-treatment variables to share information through the shared-atoms of the Bayesian nonparametric prior enables the accurate identification of principal strata while relaxing the standard assumptions of the principal stratification causal framework. Our proposed model data-adaptively identifies the principal strata and imputes missing post-treatment variables and outcomes without the need to set predefined (and subjective) thresholds on the continuous post-treatment values to identify the principal strata. Moreover, the proposed model allows for the quantification of the uncertainty in membership in the principal strata. Theorem 1 provides identification for the proposed causal estimands, while Theorem 2 illustrates how the probability of a unit belonging to a dissociative stratum is non-zero and depends on its observed covariates.
The performance of the proposed model is evaluated through an extensive set of Monte Carlo simulations where the performance of our proposed model is compared and contrasted with the performance of the proposed models by Schwartz et al. (2011) and Lu et al. (2023). The results demonstrate that our model consistently outperforms these alternatives, achieving lower bias, better principal strata identification, and accurate estimation of principal causal effects across various scenarios.
@@ 6 Discussion
The proposed model is used to assess the effectiveness of a previous revision of NAAQS in reducing levels of pm2.5and, in turn, reducing the mortality rate in treated counties. Specifically, we find significant effects on the reduction of mortality rate in areas where pollution was effectively reduced by the non-attainment designation (associative negative effects). In contrast, those areas where pollution was not measurably affected (dissociative effect) or where it actually increased (associative positive effects) did not show any significant effect on the mortality rate.